OpenStudy (faman39):

Match each actual value with the predicted value that results in a 15% error. a. Predicted : 119 b.Predicted : 85 c. Predicted : 102 d. Predicted :92 Actual :80 Actual :120 Actual :100 Actual : 140

OpenStudy (anonymous):

This looks more like it. Do you know how to start?

hero (hero):

This is just a percent change problem

hero (hero):

Percent Change Formula: \[\% \Delta = \frac{|a - b|}{a} \times 100\] a = 1st number b = 2nd number